Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day12


Crew # Log Book for January 12, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status:

Good.


Time Departed/Returned from EVA:

EVA #23: ~10:20 / ~11:30. Judah, Tonya, Nathan, Amanda, Jim, Lucinda.

EVA #24: 13:21 / 14:07. Amanda, Jim.

EVA #25: 14:22 / 14:55. Jim, Sandstorm.

EVA #26: 16:00 / 17:44. Jim, Judah.


Brief Narrative of Field Mission Results:

EVAs today consisted of all crew-member EVA Mobility Study, Water Gully analysis, Rover testing, and Satellite data comparison collection.


EVA Data/Interpretations:

EVA #23: All crew-member participation in EVA Mobility Study including testing the suits through physical activities such as football throwing, archery, and a soccer match.

EVA #24: Water Gullies were analyzed.

EVA #25: Final test of outside router and range with Sandstorm Rover. Test successful.

EVA #26: Four panoramas were taken for Satellite image comparison.


Engineering/Hab Maintenance:

Water line from outside potable tank to inside potable tank was thawed, and dip removed to reduce potential for freezing in the future.

Equipment for the Usability Engineer's experiment was repaired.


Report Transmission Schedule:

CMDR Check-in, ENGR, CMDR Report, Photos, 4 EVA Reports, 4 Science Reports


Plans for Tomorrow:

- Cooking/Cleaning - Tonya/Judah

- 07:30 - Breakfast

- 8:00 - Exercise

- Engineering checks

- EVA #27: Satellite

- 11:30 Lunch

- EVA #28: Varnish/Gullies

- EVA #26: Rover testing

- 17:30 - Dinner

- Psychophysiological crew analysis

- Reports, Food Study

- Work on experiments

- Showers - Nathan/Judah


Inventory:

Alpine Aire dehydrated food arrived today.


Support Requested:

n/a


Miscellaneous:

n/a


Commander's Report

Judah Epstein Reporting

Today on Mars was a full day of EVAs and scientific experiments.

After breakfast and crew exercise, all crew-members participated in an EVA Mobility Study whereas the Marsonauts tested their abilities of throwing a football, archery, and a game of soccer within the constraints of a spacesuit. Crew-members further realized the constraints and limitations derived from working within a spacesuit.

After lunch was several additional EVAs. The Crew Geologist and Human Factors Researcher took further analysis of water gullies. The Sandstorm Rover was further tested of the range of signal and battery usage. And additional panoramas were taken in a new geographic area for comparison with Satellite photos.

Inside the HAB, today was extended psychophysiological analysis on each crew-member. Testing on water quality of the slow sand filters was continued. And engineering repaired the water line from the outside potable tank to the inside potable tank to remove the need for "bucket brigade" and reduce potential for freezing of the line in the future. Furthermore, outreach was conducted whereas Crew 97 participated in answering questions to primary school students in England.

Tomorrow will consist of another full day of experiments and EVAs as Crew 97 is nearing completion of the Simulation.

Science Report 1 

Judah Epstein Reporting

The purpose of this project is to explore the effectiveness of a slow sand sedimentary water filtration system in a planetary environment. To simulate a planetary environment the experiment is conducted at the Mars Desert Research Station, a self-sustaining planetary research experiment and simulation expedition.

Background information regarding the project abstract, purpose, goals, and specific experiment details and information are included in the Science Report from January 6.

Data collected from Jan 11 and prior is available in this spreadsheet.

Updates on results:

Filter 1 (Martian Simulant) and Filter 2 (Commercial Sand) are not achieving optimal results as the filter seems clogged with extremely low hydraulic conductivity. Therefore not enough effluent water was produced to take measurements with the Qanta probe (Temp, SpC, DO, pH, TDS, DO%, ORP).

Turbidity results: Turbidity of effluent from all three slow sand filters has generally been increasing. Turbidity of effluent from the wetland, greywater, and HAB potable are remaining generally consistent.

Bacteria results were measured and seem to mostly be positive, but the indicators were very faint. Therefore the very faint indicators are quite possibly an indication of negative bacteria results.

On Jan 11, the Martian Regolith filter had a very faint positive indicator. The Commercial Sand filter had a positive indicator. The local MDRS sands filter had an extremely faint positive indicator. The MDRS Wetland Filtration had a very faint positive indicator. And surprisingly, the Greywater had an extremely faint positive indicator.

Science Report 2

Tonya Thompson Reporting

I performed 3 psychophysiological reading sessions on each crewmember today. Each session lasted 3 minutes.

The crewmember entered my stateroom, sat in a chair, placed finger sensors on their 3 middle fingers, and watched a white image projected onto a wall for 3 minutes. During that time, I recorded their heart rate variability (hrv) and skin conductance level (scl). The data for each session was saved to the hard drive of a desktop pc in my stateroom.

I also had each crewmember fill out a Cohen’s Perceived Stress Scale questionnaire at the end of the evening after dinner.

Science Report 3

Lucinda Land Reporting

After having issues yesterday with the computer, we decided that today we would connect the router to the computer using an ethernet cable. That proved helpful as Maxnet instantly connected to Sandstorm. We connected inside the Hab and then moved both the rover and the computer/router system outside.

Jim took the rover on the same path as with the router on the inside of the Hab. This time though the router was situated outside of the Hab. The router logged a total of 241.40 meters with the router on the outside of the Hab. This was about a 35 meter increase in distance over using the router inside the Hab. Not a significant increase in distance.

The biggest issue in using the rover is battery life. If we could have batteries that lasted longer on the rover we could take her farther. That was proven on the EVA when Jim carried Sandstorm out to the farthest point that she had been to and took her farther. Unfortunately her batteries are her limitation. We did not fully test her range. Jim could have carried her to the farthest point that her batteries lasted to see how much farther outward she could go. But other issues impede her progress such as the hilly terrain around the Hab. When she is not in sight of the router the signal diminishes and either her movement slows or stops.

Video feed was intermittent, but for the most part was not functioning. In short range Sandstorm's video works great. It was always kept at the second to lowest level of resolution.

Tomorrow I will taker her out to take snapshots of the local geology as well as take Gigapan images


Engineering Report

Nathan Wong Reporting

Generator/Electricity:

Kitty(Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): Power went out overnight.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.50

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.49

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 3 Gas cans full

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 3

Honey Oil Quantity (# of Quarts in storage): 2

ATV Oil Quantity (# of Quarts in storage): 2

Car Oil Quantity (# of Quarts in storage): none


Water Status:

Outside Potable Water Tank Level (inches from bottom): 25%

Trailer Potable Water Tank Level (inches from bottom): 75%

Hab Potable Water Tank Level (inches from bottom): empty

Potable Water Meter Reading: 40770.0

In to GreenHab Meter Reading:

Notes/Comments: Water line worked today.


GreenHab:

Condition of plants in Tank 1: None

Condition of plants in Tank 2: 3 plants

Condition of plants in Tank 3: 7 plants

Greenhab Notes/Comments: Green hab filter tank is slow to filter. Can we remove sand to make it faster?


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments: Observatory not used


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: good
Odometer at end of day: 123085
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:

ATV 2 (Spirit):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 4 (Viking II):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): off
Actual temperature upstairs (Farenheit): 73

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 57


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All webcams are now functioning properly.


General Engineering Notes/Comments:

Water line to Hab was thawed out, and dip was removed to decrease potential for freezing in the future

Repaired equipment for Tonya's experiment

Green hab filter tank is slow to filter. Can we remove sand to make it faster?


EVA 23 Report

Judah Epstein Reporting

Time: ~10:20 / ~11:30

Crew Members: Judah, Tonya, Nathan, Amanda, Jim, Lucinda.

Site Location: Next to Hab (Southern side)

Transit Mode: Foot

Objectives: EVA Mobility Study to test physical constraints in spacesuit while practicing physical activities including throwing a football, archery, and a game of soccer.

Results: Crewmembers learned about the limitations and constraints of physical activity and manual dexterity while in a spacesuit.

Lessons Learned:

Cold-weather hat lowered over crew-member's eyes and then could not see. Hats need to be smaller or properly contained.

Earpiece can easily fall off; a better system would be optimal to keep the earpiece in place.

Helmet moves a lot and can bump face during physical activity.

Helmet and spacesuit cause lowered situational awareness and visibility issues.

EVA 24 Report

Amanda Damptz Reporting

Time: 1:21 pm to 2:07 pm

Crew Members: Jim Crowell and Amanda Damptz

Site Location: East of the Hab

Transit Mode: Foot

Objectives: Take photographs of gullies.

Results: Four photographs of gullies taken. GPS coordinates taken.

Lesson Learned: The Slopes are really hard to walk up due to the saturation of the sediment after the recent snow melt. Use caution when working with equipment.

EVA 25 Report

Lucinda Land Reporting

Time: 14:22 / 14:55

Crew Members: Crewmember, Sandstorm

Site Location: Outside of Hab door, East

Transit Mode: by foot and rover

Objectives: To test range of outside router with Sandstorm

Results: Sandstorm and a crew member on EVA for 241.40 meters with the router outside the Hab. We had the router hard wired to the computer after the wi-fi failed to be seen on the computer. Maxnet connected successfully to the rover. The crew member and the rover took the route as he had done the first day when the router was inside the Hab.

Lessons Learned: If the batteries on the rover could last longer, it could probably go farther. Terrain-wise, the rover needs to be on relatively flat ground and in the line of site of the router.

EVA 26 Report

Lucinda Land Reporting

Time: 16:00 / 17:44

Crew Members: Jim Crowell, Judah Epstein

Site Location: The ATV road along the ridge

Transit Mode: ATVs

Objectives: To take panoramas along the ridge for Jim's experiment

Results: Four panoramas were taken.

Lessons Learned: EVAs for Jim's experiment need to be taken midday. The low sun casts too many shadows.